Your brother, who is prone to bearing substantial risk, suggests that you buy a security for $10,000 that promises to pay you $100,000 at the end of 15 years. What is the implied annual return or yield on this investment?
Answer #1

Let the return be I% per year

t = 15 yrs

F = P *(1+I)^t

100000 = 10000 * (1+i)^15

(1+i)^15 = 100000 / 10000 = 10

(1+i) = 10 ^(1/15) = 1.1659144

I = 1.1659144 - 1 = 0.1659144 = 16.59%